assume that y-varies inversely as x. if y=9 when x=8 find y when x=6
Answer by stanbon(75887)

assume that y-varies inversely as x.
y = k/x
------
Solve for "k" using "y=9 when x=8":
9 = k/8
k = 72
----
find y when x=6
Equation:
y = 72/x
y = 72/6
y = 12
===========
Cheers,
Stan H.
------